OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\log_{1 / 2 } 8\]

OpenStudy (anonymous):

log( base 1/2) 8 = x <=> (1/2)^x = 8 1/(2^x) = 8 2^x = 1/8 1/(2^(-x)) = 1/8 2^(-x) = 8 log(base 2)2^(-x) = log(base 2)8 -x*log(base 2)2 = 3 (because 2^3 = 2*2*2 = 8) -x = 3 (because log(base 2)2 =1, because 2^1=2) x = -3

OpenStudy (anonymous):

Same thing 1/2^x = 8 x = -3

OpenStudy (anonymous):

so the x becomes the exponent thank you!! :-) @MathFan12 @kevinzhengli